OpenStudy (anonymous):

Suppose Q and R are independent events. Find P (Q and R) P(Q)=2/5, P(R)=3/7

OpenStudy (anonymous):

if they are independent, then \[P(Q\cap R)=P(Q)\times P(R)\]

OpenStudy (anonymous):

in other words, multiply the two probabilities

OpenStudy (anonymous):

I tried and keep getting .92

OpenStudy (anonymous):

\[\frac{2}{5} \times \frac{3}{7} \ne 0.92\]...

OpenStudy (anonymous):

\[\frac{2}{5}\times \frac{3}{7}=\frac{6}{35}\]

OpenStudy (anonymous):

Oh, thank you :D Would you mind helping me on one more problem?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

P(Q)=1/3, P(R)=1/10